40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 24 11:30:27 2011, jamie, jenne, kiwamu, suresh, steve, Update, SUS, ETMX wiped, replaced, door on 
    Reply  Thu Aug 25 08:12:09 2011, steve, Update, SUS, ETMY & ITMY wiped, replaced, vac door on 
       Reply  Thu Aug 25 13:10:42 2011, Jenne, Update, SUS, Drag wiping 
Message ID: 5300     Entry time: Thu Aug 25 08:12:09 2011     In reply to: 5295     Reply to this: 5301
Author: steve 
Type: Update 
Category: SUS 
Subject: ETMY & ITMY wiped, replaced, vac door on 

Quote:

We've closed up ETMX:

  • the optic was drag wiped
  • the suspension tower was put back in place
  • earthquake stops were backed off the appropriate number of turns, and de-ionized
  • chamber door was put on

 jamie, jenne, kiwamu, suresh, steve

 

ETMY and ITMY were treated the same way as ETMX. The BS chamber was closed with heavy vac door yesterday also. The IOO access connector's inner jamnuts are torqued to 45 ft/lbs as all vac door bolts.

The vac envelope is ready for pumpdown condition, except ITMX chamber with light atm door cover.

Jenne will summeries the condition of dust on the  TMs before and after the drag wipes.

ELOG V3.1.3-